if 2cos2θ=√3, find the value of  θ ? Solution Please?

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
cos 2 theta = cos (pi/6)
 
2 theta = 2*n*pi ± pi/6
 
theta = n *pi ± pi/12
 
now put values of n and find different values of theta